A student planned an experiment that would use 0.10 M propionic acid, HC3H5O2. Calculate the value of [H+] and the pH for this solution. For propionic acid, Ka = 1.3
Answer #1

HC3H5O2 <--> H+ and C3H5O2-

Ka = [H+][A-] /[HA]

Ka = 1.3*10^-5

[H+] =[A-] = x

[HA] = 0.1

1.3*10^-5 = x*x /(0.1)

1.3*10^-6 = X^2

x = sqrt(1.3*10^-6 ) = 0.00114

[H+] = 0.00114 M

pH = -log(H+) = -log(0.00114) = 2.94